The daily newspaper always follows a particular format. Each even-numbered page contains 6 articles and each odd-numbered page contains 7 articles. If today paper has 36 pages, how many articles does it contain ?
(A) 282 (B) 249 (C) 234 (D) 220 (E) 180

Answer by checkley77(12844)

THERE ARE 36/2=18 PAIR OF ODD-EVEN PAGES.
EACH PAIR HAS 6+7=13 ARTICLES.
18*13=234 TOTAL ARTICLES.
ANSWER C).